# ViewSRD: 3D Visual Grounding via Structured Multi-View Decomposition

### v1

@@ 4.5 Ablation Study
1.0 1.0 1.0 70.5
Analysis of Loss Weight. Table 7 presents an ablation study on the weighting scheme of the loss functions. The best performance 70.5% is achieved when equal weights are assigned to $\mathcal{L}_{Object}$, $\mathcal{L}_{Sent}$, and $\mathcal{L}_{Ref}$, validating our multi-loss design.
5 Conclusion
A limitation of ViewSRD is its assumption that complex queries can be fully decomposed without overlapping relationships. While decomposition into overlapping relations does not degrade performance, it diminishes the intended benefits of simplification. Future work will explore adaptive query representations to better preserve contextual dependencies.

### v2

@@ Appendix E Loss Function
1.0 1.0 1.0 69.8
Table 7 presents an ablation study on the weighting scheme of the loss components $\mathcal{L}_{Ref}$, $\mathcal{L}_{Obj}$, and $\mathcal{L}_{Sent}$. The best accuracy of 69.9% is achieved when the loss weights are set to $\lambda_{Obj}=1.0$, $\lambda_{Ref}=0.5$, and $\lambda_{Sent}=0.5$, confirming the effectiveness of our multi-loss formulation with this specific weighting scheme.
Notably, omitting either $\mathcal{L}_{Obj}$ or $\mathcal{L}_{Sent}$ leads to significant performance degradation (to 63.9% and 69.1%, respectively), highlighting the importance of jointly modeling object-level and sentence-level supervision. We also observe that moderate weighting (e.g., 0.25 or 0.5) for $\mathcal{L}_{Obj}$ and $\mathcal{L}_{Sent}$ improves over their absence, but still underperforms compared to the balanced setting. These results validate that all three components contribute synergistically to grounding performance.